Show that ∞n=0 (x^n)/(n!)converges for all x ∈ R.


1
Expert's answer
2021-02-24T06:40:00-0500

Using the ration test, the radius of convergence of the series "\displaystyle \sum_{n=0}^{+\infty}\frac{x^n}{n!}" is

"r=\lim_{n\rightarrow\infty}\left|\frac{c_n}{c_{n+1}}\right|=\lim_{n\rightarrow\infty}\left|\frac{\frac{1}{n!}}{\frac{1}{(n+1)!}}\right|=\lim_{n\rightarrow\infty}n+1=+\infty."


That is, for all "x\in\mathbb{R}", the series is convergent.